OpenStudy (moongazer):

16 years from now Peter will be twice as old as Paul. The current sum of the ages of Peter and Paul is 4. How old is Paul right now? Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (anonymous):

let the present age of peter be x and paul be y x+y=4------eqn1 after 16 years peter's age=x+16 paul's age=y+16 x+16=2(y+16) x+16=2y+32------eqn2 now solve for x and y

OpenStudy (moongazer):

can you make a table like solution? because my teacher is doing that and it is easier to understand